A Rod of Mass m and length l is rotating with angular speed w about an axis, passing through it's end point and perpendicular to it's length. Find the tension in the string at distance x from the end

Grade:11

2 Answers

Arun
4 years ago
As no external torque is present, there is no angular acceleration and thus angular velocity remains constant. Also as the angular velocity is constant the kinetic energy too remains conserved. Also by definition, for no external torque the angular momentum about the center of mass remains conserved
Vikas TU
4 years ago
 mass of the element = dm =m/l dx 
The force on the element towards the axis =T−(T+dT)=−dT
−dT=(dm)ω^2x = (m/ldx)w^2x 
T = -mw^2/l * x^2/2 + constant 
x = 1 , T = 0 
constant  1/2 * (mw^2l^2)/l
T = 1/2 mw^2/l (l^2-x^2)
Good Luck